# Learning Priors for Non Rigid SfM from Casual Videos

### v1

@@ 6 Conclusion
We presented TracksTo4D, a novel deep-learning framework that directly translates 2D motion tracks from casual videos into their corresponding non-rigid structure and motion. Our approach features a deep learning architecture that considers the symmetries in the problem with designed intrinsic constraints to handle the ill-posed nature of this problem. Notably, our network was trained using only raw supervision of 2D point tracks extracted by an off-the-shelf method [18] without any 3D supervision. Yet, it implicitly learned to predict camera poses and non-rigid structures while identifying the non-rigid parts. During inference, our method demonstrates faster processing times compared to previous methods while achieving comparable accuracy. Furthermore, our network exhibits strong generalization capabilities, performing well even on semantic categories that were not present in the training data.
@@ Acknowledgments
Appendix 0.A Supplementary Results
0.A.1 Video Results
@@
0.A.2 Extended Quantitative Evaluation
Per-sequence and mean quantitative comparisons for our 21 pet test videos are presented in Tab. 5 and Tab. 6. We also show per-sequence fine-tuned results of 2000 iterations starting from our final checkpoint. The fine-tuning is done as a post-processing by minimizing the original loss function on the given test video. This takes about $10$ minutes per case which is still fast compared to the run time of the previous methods that predicted dynamic depths (see Tab. 1).
0.A.3 Generalization Beyond Pet Videos
We further applied our pretrained network on the MPI Sintel dataset [9] for the evaluation set defined by [64] of $14$ sequences. We note that the pet domain that was used for training our network is very different from the Sintel domain which contains fast motion of multiple characters. Yet, we can see that due to the strong prior collected by our network, it generalizes to some extent to the domain of Sintel. Fig. 6 shows a qualitative evaluation of our movement level predictions ($\gamma$). As can be seen, despite the domain gap (compared to Fig. 5), our network identifies the non-rigid movement in the videos. We further evaluated the camera pose accuracy in Tab. 7. Even though our mean accuracy is not the best among all methods, our method gets the most accurate results on some metrics for several sequences. Training on a video dataset in a similar domain to Sintel could possibly improve the generalization of our method for Sintel. Unfortunately, the Sintel dataset contains only 23 sequences.
Figure 6: $\gamma$ Visualization for Sintel. We show a visualization of the $\gamma$ outputs of our network. In each video sequence, we show the input tracks, where each color visualizes its movement level value, $\gamma$. Purple marks rigid points with low $\gamma$ whereas red marks non-rigid points with high $\gamma$. Note that our network was only trained on pet videos, and did not see any videos in the domain of Sintel. Yet, at inference time with no further optimization, it identifies the non-rigid movement successfully.
Appendix 0.B Implementation Details
@@
0.B.1 Architecture
Here we provide a precise definition for our equivariant layer: $F:\mathbb{R}^{N\times P\times d}\rightarrow\mathbb{R}^{N\times P\times d}$. The computation of $F$ is split into four steps, as described below:
$\mathbf{\bar{q}}_{ij}=\bar{W}^{Q}M_{ij},~{}\mathbf{\bar{k}}_{ij}=\bar{W}^{K}M_{ij},~{}\mathbf{\bar{v}}_{ij}=\bar{W}^{V}M_{ij},$ (12a)
$\bar{M}_{ij}=\sum_{i^{\prime}=1}^{N}\frac{\exp(\mathbf{\bar{q}}_{ij}\cdot\mathbf{\bar{k}}_{i^{\prime}j})}{\sum_{l=1}^{N}\exp(\mathbf{\bar{q}}_{ij}\cdot\mathbf{\bar{k}}_{lj})}\mathbf{\bar{v}}_{i^{\prime}j}$ (12b)
$\mathbf{q}_{ij}=W^{Q}\bar{M}_{ij},~{}\mathbf{k}_{ij}=W^{K}\bar{M}_{ij},~{}\mathbf{v}_{ij}=W^{V}\bar{M}_{ij},$ (12c)
$F(M)_{ij}=\sum_{j^{\prime}=1}^{P}\frac{\exp(\mathbf{q}_{ij}\cdot\mathbf{k}_{ij^{\prime}})}{\sum_{l=1}^{P}\exp(\mathbf{q}_{ij}\cdot\mathbf{k}_{il})}\mathbf{v}_{ij^{\prime}}$ (12d)
Here, $M_{i,j}\in\mathbb{R}^{d}$ are the features associated with the $j$-th point in the $i$-th frame. The attention mechanism defined in the first two equations above (with additional temporal positional encoding) replaces the translation equivariant function $L$ applied to the columns of $M$ (Eq.(3)), and the attention in the last two equations above implements the point cloud aggregation (summation) (also in Eq.(3)) applied to the rows of $M$. As commonly done we use 16 attention head transformations [46]. We add standard temporal positional encoding for the transformers with respect to the $N$ (time) dimension as described in [46]. In our implementation, we added positional encoding only in the first layer.
Architecture Technical Details
For learning high frequencies we map each input coordinate to sinusoidal functions as in [28] with $L=12$. We use 3 pairs of attention layers, each of frames attention followed by point attention. Each point (after sinusoidal functions embedding) is mapped into $\mathbb{R}^{256}$ with a linear layer. Each attention layer is a function of the form $F:\mathbb{R}^{N\times P\times 256}\rightarrow\mathbb{R}^{N\times P\times 256}$ (see details above). Each attention uses 16 heads with $K,Q,V\in\mathbb{R}^{(N\text{ or }P)\times 64}$ followed by a fully connected network with 1 hidden layer of $2048$ features. We then average over the rows to get per-point features $P_{0}\in\mathbb{R}^{P\times 256}$ and over the columns to get per-frame features $F_{0}\in\mathbb{R}^{N\times 256}$. Finally, we map $P_{0}$ to per-point outputs $P_{1}\in\mathbb{R}^{P\times(3K+1)}$ ( K basis points and $\gamma$) with a linear layer, and $F_{0}$ into per-camera outputs $F_{1}\in\mathbb{R}^{N\times(6+3+K)}$ ($6$ for the rotation parameters [67], $3$ for the camera center, and K coefficients) using a conv layer with a kernel size of $31$.
0.B.2 Other Implementation Details
@@
For building $M\in\mathbb{R}^{N\times P\times 3}$ we use the implementation of [18]. We sample a uniform grid of $15\times 15$ 2D points, starting from frame number $0,20,40,\dots$, and then track these points throughout the entire video (backward and forward). In Tab. 4 we show the effect of reducing the grid size at inference time, in terms of camera pose accuracy and running time. During training, at each iteration, we randomly sample 20-50 frames from the training videos and 100 point tracks, i.e. $20\leq N\leq 50$ and $P=100$. When sampling cameras and point tracks of size $N\times P$ from a larger matrix of size $N^{\prime}\times P^{\prime}$ we only take a point track if its starting tracking time is in the range $[t-\frac{N}{2},t+\frac{3N}{2}]$, where t is the first sampled index. At inference time we take all the available point tracks. In both, training and inference time, we keep only point tracks that are observed in more than $10$ frames.
@@
We used the RGBD camera of the iPhone 11 to record our 21 test videos of dogs and cats. Each frame has a resolution of $640\times 480$ pixels. Note that the training set contained various types of resolutions. All pet owners who were photographed gave their permission for the animals to be photographed. For evaluation only, we define a point track as dynamic if its associated GT mask value is $1$ for at least $40$ frames. The GT masks are obtained by running [68] and searching for labels of dogs and cats. They were only used for evaluation and not used by our method at all.
@@
After inference, as optional refinement, we take the output rigid approximation $B_{1}\in\mathbb{R}^{P\times 3}$ and the output camera poses $\{R_{1},\dots,R_{N}\}$, $\{\mathbf{t}_{1},\dots,\mathbf{t}_{N}\}$ and apply Bundle Adjustment (BA). We use a 3D world point from $B_{1}$ if its associated $\gamma$ is below $0.008$. We optimize reprojection errors of a given observation $M_{i,j}$, only if it is observed, i.e. $M_{i,j}^{o}=1$ and if the initial reprojection error is below $10$ pixels. We use the BA implementation provided by [29], which is based on the Ceres package [2].
@@
In all training setups, we used: $\lambda_{\text{Reprojection}}=50.0$, $\lambda_{\text{Rigid}}=1.0$, $\lambda_{\text{Negative}}=1.0$, $\lambda_{\text{Sparse}}=0.001$. At the beginning of the training, we pre-train the camera poses to be located behind and facing the origin. This prevents cases in which the cameras are located in the middle of the initial point cloud s.t. many points have negative depths, which may result in bad convergence. More specifically, the pre-train loss is:
$\mathcal{L}_{\text{Pretrain}}=\frac{1}{N}\sum_{i=1}^{N}\frac{1}{100}\left\lVert\mathbf{t}_{i}-[0,0,-15]^{T}\right\rVert^{2}+\left\lVert R_{i}-I\right\rVert_{F}^{2}$
@@
Table 5: Depth Accuracy. We show a comparison to previous methods on the predicted depth for the point trajectories compared to their GT depths. We compare 4 ways of running our method. Ours (C&D), Ours (C): Our inference time outputs for a model that was trained on cats and dogs or only on cats respectively. Ours (C&D) FT, Ours (C) FT: The outputs of our model trained on cats and dogs or cats respectively, after fine-tuning our losses for each specific video. As can be seen, fine-tuning improves our accuracy even more.
@@
$\delta<1.25\uparrow$ Dyn 0.45 0.95 0.92 0.84 0.89 0.90 0.91
All 0.65 0.80 0.97 0.86 0.96 0.91 0.96
$\delta<1.25^{2}\uparrow$ Dyn 0.68 1.00 0.99 0.98 0.99 0.98 0.99
@@
All 0.96 0.93 1.00 1.00 0.99 1.00 0.99
Table 6: Camera Poses Accuracy. We show a comparison to previous methods on the predicted camera poses. We compare 3 ways of running our method. Ours (C): Our inference time outputs (total inference time of 0.16 seconds) for a model that was trained only on cats. Ours (C)+BA: Our inference time outputs, followed by a short Bundle Adjustment (total inference time of 0.4 seconds) for a model that was trained only on cats. Ours (C) FT: The outputs of the model that was trained only on cats after fine-tuning our losses for each specific video (total running time of about 10 minutes). As can be seen, after BA, our results are the most accurate compared to the other methods, and fine-tuning improves our accuracy even more.
@@
Table 7: Camera Poses Accuracy on Sintel. We show a comparison to previous methods on the predicted camera poses. We compare 4 ways of running our method. Ours (C), Ours (C&D): Our inference time outputs for a model that was trained only on cats or on cats and dogs respectively. Ours (C)+BA, Ours (C&D)+BA: The outputs of the model trained only on cats or on cats and dogs respectively, after further applying short BA refinement. The numbers for [21],[43] and [64] were obtained from [64].
RCVD[21] DROID- Particle CasualSAM[62] Ours Ours Ours Ours
SLAM[43] SfM[64] (C) (C)+BA (C&D) (C&D)+BA
alley_2 ATE(mm) 0.026 0.057 0.002 0.0008 0.043 0.027 0.055 0.046
RPE T.(mm) 0.056 0.035 0.0007 0.0004 0.033 0.040 0.034 0.041
RPE R.(deg.) 0.821 1.047 0.009 0.008 0.468 0.507 0.511 0.531
ambush_4 ATE(mm) 0.171 0.104 0.068 0.011 0.050 0.021 0.055 0.022
RPE T.(mm) 0.048 0.035 0.027 0.007 0.026 0.026 0.026 0.026
RPE R.(deg.) 3.025 1.385 0.473 0.079 0.398 0.319 0.382 0.324
ambush_5 ATE(mm) 0.230 0.112 0.002 0.008 0.057 0.010 0.035 0.009
RPE T.(mm) 0.046 0.029 0.001 0.003 0.018 0.008 0.015 0.006
RPE R.(deg.) 4.105 1.580 0.055 0.084 0.430 0.228 0.338 0.149
ambush_6 ATE(mm) 0.199 0.289 0.269 0.042 0.241 0.223 0.191 0.141
RPE T.(mm) 0.112 0.078 0.081 0.024 0.088 0.129 0.072 0.078
RPE R.(deg.) 4.147 4.596 0.951 0.482 1.928 3.364 1.602 1.849
cave_2 ATE(mm) 0.596 0.351 0.961 0.967 0.217 0.217 0.460 0.460
RPE T.(mm) 0.171 0.172 0.142 0.170 0.094 0.094 0.113 0.113
RPE R.(deg.) 7.508 5.489 3.678 2.674 1.781 1.781 2.634 2.634
cave_4 ATE(mm) 0.179 0.155 0.068 0.009 0.055 0.006 0.115 0.025
RPE T.(mm) 0.087 0.035 0.012 0.003 0.024 0.006 0.028 0.021
RPE R.(deg.) 2.040 2.710 0.409 0.052 0.244 0.044 0.294 0.171
market_2 ATE(mm) 0.032 0.011 0.003 0.001 0.027 0.002 0.033 0.003
RPE T.(mm) 0.018 0.006 0.010 0.0006 0.004 0.002 0.007 0.002
RPE R.(deg.) 0.141 0.036 0.041 0.010 0.107 0.008 0.091 0.008
market_5 ATE(mm) 1.213 0.912 0.012 0.005 0.891 1.187 0.859 0.525
RPE T.(mm) 0.762 0.293 0.006 0.002 0.212 0.557 0.209 0.395
RPE R.(deg.) 1.863 3.334 0.034 0.016 1.252 6.946 1.489 3.285
market_6 ATE(mm) 0.248 0.057 0.018 0.008 0.187 0.163 0.220 0.152
RPE T.(mm) 0.214 0.037 0.007 0.002 0.103 0.097 0.095 0.086
RPE R.(deg.) 0.817 1.296 0.120 0.019 1.003 0.850 0.834 0.695
shaman_3 ATE(mm) 0.054 0.001 0.0005 0.002 0.002 0.002 0.003 0.002
RPE T.(mm) 0.023 0.002 0.0004 0.0004 0.0007 0.0004 0.001 0.0004
RPE R.(deg.) 0.718 0.199 0.072 0.036 0.060 0.042 0.087 0.039
sleeping_1 ATE(mm) 0.029 0.011 0.008 0.008 0.002 0.002 0.002 0.002
RPE T.(mm) 0.019 0.006 0.001 0.001 0.001 0.0006 0.001 0.0006
RPE R.(deg.) 0.668 0.479 0.042 0.040 0.070 0.040 0.096 0.040
sleeping_2 ATE(mm) 0.043 0.005 0.0007 0.0002 0.005 0.0006 0.005 0.0006
RPE T.(mm) 0.049 0.018 0.0002 0.0002 0.003 0.0007 0.003 0.0007
RPE R.(deg.) 0.446 0.139 0.006 0.008 0.108 0.018 0.088 0.018
temple_2 ATE(mm) 1.245 0.073 0.011 0.086 0.189 0.011 0.217 0.009
RPE T.(mm) 0.394 0.348 0.002 0.008 0.029 0.004 0.032 0.003
RPE R.(deg.) 1.318 1.298 0.019 0.131 0.287 0.019 0.340 0.018
temple_3 ATE(mm) 0.769 0.310 0.381 0.587 0.350 0.510 0.329 0.473
RPE T.(mm) 0.161 0.093 0.149 0.136 0.114 0.190 0.114 0.138
RPE R.(deg.) 20.592 3.230 1.583 0.109 1.672 2.110 1.458 1.176
Mean ATE(mm) 0.360 0.175 0.129 0.124 0.165 0.170 0.184 0.133
RPE T.(mm) 0.154 0.085 0.031 0.025 0.054 0.082 0.054 0.065
RPE R.(deg.) 3.444 1.916 0.535 0.268 0.701 1.163 0.732 0.781

### v2

@@ 5 Conclusions and limitations
We presented TracksTo4D, a novel deep-learning framework that directly maps 2D motion tracks from casual videos into their corresponding dynamic structure and camera motion. Our approach features a deep learning architecture that considers the symmetries in the problem with designed intrinsic constraints to handle the ill-posed nature of this problem. Notably, our network was trained using only raw supervision of 2D point tracks extracted by an off-the-shelf method [18] without any 3D supervision. Yet, it implicitly learned to predict camera poses and 3D structures while identifying the dynamic parts. During inference, our method demonstrates significantly faster processing times compared to previous methods while achieving comparable accuracy. Furthermore, our network exhibits strong generalization capabilities, performing well even on semantic categories that were not present in the training data.
Limitations and future work.
While our experiments demonstrated that our network is efficient, accurate, and capable of generalizing to unseen video categories, there are several limitations and future work directions that we would like to address. First, Our method cannot handle videos with too rapid motion, and in general, is limited by the accuracy of the tracking method [18]. We note that any future improvements with point tracking in terms of accuracy and inference time will directly improve our method as well. Our method assumes enough motion parallax to constrain the depth values and fails to generate accurate camera poses without it. A future and interesting work would be to try combining depth-from-single-image prior as additional inputs to our network for handling cases with minimal motion parallax and improving accuracies.
@@ Acknowledgments
Appendix A Supplementary results
A.1 Video results
@@
A.2 Extended quantitative evaluation
Per-sequence and mean quantitative comparisons for our 21 pet test videos are presented in Tab. LABEL:tab::depth_per_seq and Tab. LABEL:tab::cam_per_seq. Tables with similar structure for the out-of-domain dataset are presented in Tab. LABEL:tab::depth_per_seq_nvidia and Tab. LABEL:tab::cam_per_seq_nvidia.
Appendix B Implementation details
@@
B.1 Architecture technical details
For learning high frequencies we map each input coordinate to sinusoidal functions as in [28] with $L=12$. We use 3 pairs of attention layers, each of frames attention followed by point attention. Each point (after sinusoidal functions embedding) is mapped into $\mathbb{R}^{256}$ with a linear layer. Each attention layer is a function of the form $F:\mathbb{R}^{N\times P\times 256}\rightarrow\mathbb{R}^{N\times P\times 256}$ (see details above). Each attention uses 16 heads with $K,Q,V\in\mathbb{R}^{(N\text{ or }P)\times 64}$ followed by a fully connected network with 1 hidden layer of $2048$ features. We then average over the rows to get per-point features $P_{0}\in\mathbb{R}^{P\times 256}$ and over the columns to get per-frame features $F_{0}\in\mathbb{R}^{N\times 256}$. Finally, we map $P_{0}$ to per-point outputs $P_{1}\in\mathbb{R}^{P\times(3K+1)}$ ( K basis points and $\gamma$) with a linear layer, and $F_{0}$ into per-camera outputs $F_{1}\in\mathbb{R}^{N\times(6+3+K-1)}$ ($6$ for the rotation parameters [64], $3$ for the camera center, and $K-1$ linear coefficients) using a convolutional layer with a kernel size of $31$.
B.2 Training details
In total, we used $733$ cat videos and $753$ dog videos for training. We trained our networks for $7000$ and $3500$ epochs for the single-class and multi-class setups respectively. Training our method lasts about one week on a single Tesla V100 GPU with 32GB memory. We used the Adam optimizer [20] with a learning rate of $10^{-4}$. Our method assumes known camera internal parameters which are provided by the dataset and used to normalize the point tracks as a preprocessing step.
B.3 Other implementation details
@@
For building $M\in\mathbb{R}^{N\times P\times 3}$ we use the implementation of [18]. We sample a uniform grid of $15\times 15$ 2D points, starting from frame number $0,20,40,\dots$, and then track these points throughout the entire video (backward and forward). In Tab. 4 we show the effect of reducing the grid size at inference time, in terms of camera pose accuracy and running time. During training, at each iteration, we randomly sample 20-50 frames from the training videos and 100 point tracks, i.e. $20\leq N\leq 50$ and $P=100$. When sampling cameras and point tracks of size $N\times P\times 3$ from a larger tensor of size $N^{\prime}\times P^{\prime}\times 3$ we only take a point track if its starting tracking time is in the range $[t-\frac{N}{2},t+\frac{3N}{2}]$, where t is the first sampled index. At inference time we take all the available point tracks. In both, training and inference time, we keep only point tracks that are observed in more than $10$ frames.
Finetuning details
For our fine-tuning (FT) in the main paper, we applied per-sequence fine-tunning of 500,100 iterations starting from our final checkpoint, for pets,out-of-domain data respectively. The fine-tuning is done as a post-processing by minimizing the original loss function on the given test video.
@@
We used the RGBD camera of the iPhone 11 to record our 21 test videos of dogs and cats. Each frame has a resolution of $640\times 480$ pixels. Note that the training set contained various types of resolutions. All pet owners who were photographed gave their permission for the animals to be photographed. For evaluation only, we define a point track as dynamic if its associated GT mask value is $1$ for at least $40$ frames. The GT masks are obtained by running [65] and searching for labels of dogs and cats. They were only used for evaluation and not used by our method at all. We verified that this data includes enough dynamic motion, by also including several videos that COLMAP failed to reconstruct without the masks. We further verified manually that the camera trajectories look reasonable.
The out-of-domain dataset contains video sequences with 24 frames, each of resolution of $546\times 288$. The GT dynamic masks are provided by the dataset. For this dataset, for evaluation only, we define a point track as dynamic if its associated GT mask value is $1$ for at least $15$ frames.
@@
After inference, as optional refinement, we take the output static approximation $B_{1}\in\mathbb{R}^{P\times 3}$ and the output camera poses $\{R_{1},\dots,R_{N}\}$, $\{\mathbf{t}_{1},\dots,\mathbf{t}_{N}\}$ and apply Bundle Adjustment (BA). We use a 3D world point from $B_{1}$ if its associated $\gamma$ is below $0.008$ for the pets dataset and $0.005$ for out-of-domain dataset. We optimize reprojection errors of a given observation $M_{i,j}$, only if it is observed, i.e. $M_{i,j}^{o}=1$ and if the initial reprojection error is below $10$ pixels. We use the BA implementation provided by [29], which is based on the Ceres package [2].
Running times
All inference running times were computed on a machine with NVIDIA RTX A6000 GPU and Intel(R) Core(TM) i7-9800X 3.80GHz CPU. Extracting point tracks with [18] took 8.6 and 2.5 seconds for each video on the pet-videos and out-of-domain videos respectively and included in the running time tables as part of our method inference time.
@@
In all training setups, we used: $\lambda_{\text{Reprojection}}=50.0$, $\lambda_{\text{Static}}=1.0$, $\lambda_{\text{Negative}}=1.0$, $\lambda_{\text{Sparse}}=0.001$. At the beginning of the training, we pre-train the camera poses to be located behind and facing the origin. This prevents cases in which the cameras are located in the middle of the initial point cloud s.t. many points have negative depths, which may result in bad convergence. More specifically, the pre-train loss is: $\mathcal{L}_{\text{Pretrain}}=\frac{1}{N}\sum_{i=1}^{N}\frac{1}{100}\left\lVert\mathbf{t}_{i}-[0,0,-15]^{T}\right\rVert^{2}+\left\lVert R_{i}-I\right\rVert_{F}^{2}$.
@@
Table 5: Depth accuracy, pet test set. We show a comparison to previous methods on the predicted depth for the point trajectories compared to their GT depths. We compare 4 ways of running our method. Ours (C&D), Ours (C): Our inference time outputs for a model that was trained on cats and dogs or only on cats respectively. Ours (C&D) FT, Ours (C) FT: The outputs of our model trained on cats and dogs or cats respectively, after fine-tuning our losses for each specific video. As can be seen, fine-tuning improves our accuracy even more.
@@
$\delta<1.25\uparrow$ Dyn 0.45 0.95 0.92 0.84 0.91 0.90 0.91
All 0.65 0.80 0.97 0.86 0.96 0.91 0.97
$\delta<1.25^{2}\uparrow$ Dyn 0.68 1.00 0.99 0.98 0.98 0.98 0.98
@@
All 0.96 0.93 1.00 1.00 1.00 1.00 1.00
[…]
All 9.39E+07 2.04E+05 0.02 0.04 0.02 0.04 0.02
$\delta<1.25\uparrow$ Dyn 0.29 0.17 0.10 0.18 0.15 0.16 0.16
All 0.10 0.11 0.03 0.10 0.03 0.11 0.03
$\delta<1.25^{2}\uparrow$ Dyn 0.26 0.05 0.02 0.03 0.01 0.03 0.01
All 0.07 0.07 0.01 0.02 0.01 0.03 0.01
$\delta<1.25^{3}\uparrow$ Dyn 0.15 0.00 0.00 0.00 0.00 0.00 0.00
All 0.05 0.05 0.01 0.01 0.00 0.01 0.00
Table 6: Camera poses accuracy for pets. We show a comparison to previous methods on the predicted camera poses. We compare 3 ways of running our method. Ours (C): Our inference time outputs (total inference time of 0.16 seconds) for a model that was trained only on cats. Ours (C)+BA: Our inference time outputs, followed by a short Bundle Adjustment (total inference time of 0.4 seconds) for a model that was trained only on cats. Ours (C) FT: The outputs of the model that was trained only on cats after fine-tuning our losses for each specific video (total running time of about 5 minutes). As can be seen, after BA, our results are the most accurate compared to the other methods, and fine-tuning improves our accuracy even more.
@@
Table 7: Depth accuracy for out-of-domain data [58]. We show a comparison to previous methods on the predicted depth for the point trajectories compared to their GT depths. We compare 2 ways of running our method. Ours (C&D): Our inference time outputs for a model that was trained on cats and dogs. Ours (C&D) FT: The outputs of our model trained on cats and dogs, after fine-tuning our losses for each specific video. As can be seen, fine-tuning improves our accuracy even more.
RCVD [22] MiDaS[4] CasualSAM[59] Ours (C&D) Ours (C&D) FT
[…]
All 0.14 0.34 0.02 0.06 0.01
$\delta<1.25\uparrow$ Dyn 0.42 0.89 0.98 0.98 0.98
All 0.62 0.72 0.99 0.99 0.99
$\delta<1.25^{2}\uparrow$ Dyn 1.00 0.98 0.99 0.99 0.99
All 1.00 0.81 1.00 1.00 1.00
$\delta<1.25^{3}\uparrow$ Dyn 1.00 0.99 1.00 1.00 1.00
All 1.00 0.88 1.00 1.00 1.00
[…]
All 0.14 4.87E+05 0.01 0.06 0.01
$\delta<1.25\uparrow$ Dyn 0.97 0.95 1.00 0.99 1.00
All 0.83 0.76 1.00 0.97 1.00
$\delta<1.25^{2}\uparrow$ Dyn 1.00 0.99 1.00 1.00 1.00
All 1.00 0.86 1.00 1.00 1.00
$\delta<1.25^{3}\uparrow$ Dyn 1.00 0.99 1.00 1.00 1.00
All 1.00 0.93 1.00 1.00 1.00
[…]
All 0.05 4.75E+04 0.01 0.06 0.01
$\delta<1.25\uparrow$ Dyn 0.94 0.03 0.99 0.98 0.98
All 0.98 0.67 1.00 1.00 1.00
$\delta<1.25^{2}\uparrow$ Dyn 1.00 0.04 1.00 0.98 1.00
All 1.00 0.82 1.00 1.00 1.00
$\delta<1.25^{3}\uparrow$ Dyn 1.00 0.18 1.00 1.00 1.00
All 1.00 0.86 1.00 1.00 1.00
[…]
All 0.12 0.59 0.02 0.05 0.04
$\delta<1.25\uparrow$ Dyn 0.77 0.07 0.99 0.96 0.96
All 0.86 0.22 0.99 0.97 0.97
$\delta<1.25^{2}\uparrow$ Dyn 1.00 0.30 1.00 1.00 0.99
All 1.00 0.38 1.00 1.00 1.00
$\delta<1.25^{3}\uparrow$ Dyn 1.00 0.75 1.00 1.00 1.00
All 1.00 0.64 1.00 1.00 1.00
[…]
All 0.30 7.67E+03 0.07 0.15 0.08
$\delta<1.25\uparrow$ Dyn 0.36 0.49 0.93 0.64 0.89
All 0.44 0.59 0.96 0.78 0.94
$\delta<1.25^{2}\uparrow$ Dyn 0.61 0.72 0.99 0.98 0.94
All 0.71 0.78 0.98 0.91 0.97
$\delta<1.25^{3}\uparrow$ Dyn 0.67 0.83 1.00 0.98 0.94
All 0.82 0.87 0.99 0.98 0.98
[…]
All 0.10 1.09 0.02 0.05 0.04
$\delta<1.25\uparrow$ Dyn 0.89 0.59 0.76 0.93 0.93
All 0.92 0.29 0.99 1.00 0.99
$\delta<1.25^{2}\uparrow$ Dyn 1.00 0.95 0.97 0.97 0.97
All 1.00 0.42 1.00 1.00 1.00
$\delta<1.25^{3}\uparrow$ Dyn 1.00 0.97 1.00 1.00 0.97
All 1.00 0.53 1.00 1.00 1.00
[…]
All 2.06E+06 0.22 0.03 0.14 0.05
$\delta<1.25\uparrow$ Dyn 0.26 0.81 0.99 0.94 1.00
All 0.50 0.71 1.00 0.79 1.00
$\delta<1.25^{2}\uparrow$ Dyn 0.99 0.99 1.00 1.00 1.00
All 0.88 0.94 1.00 1.00 1.00
$\delta<1.25^{3}\uparrow$ Dyn 1.00 1.00 1.00 1.00 1.00
All 1.00 0.99 1.00 1.00 1.00
[…]
All 0.10 1.62E+06 0.02 0.05 0.02
$\delta<1.25\uparrow$ Dyn 0.91 0.87 1.00 1.00 1.00
All 0.84 0.68 1.00 0.99 1.00
$\delta<1.25^{2}\uparrow$ Dyn 1.00 0.91 1.00 1.00 1.00
All 1.00 0.71 1.00 1.00 1.00
$\delta<1.25^{3}\uparrow$ Dyn 1.00 0.92 1.00 1.00 1.00
All 1.00 0.72 1.00 1.00 1.00
[…]
All 2.58E+05 2.70E+05 0.03 0.08 0.03
$\delta<1.25\uparrow$ Dyn 0.69 0.59 0.95 0.93 0.97
All 0.75 0.58 0.99 0.94 0.99
$\delta<1.25^{2}\uparrow$ Dyn 0.95 0.73 0.99 0.99 0.99
All 0.95 0.72 1.00 0.99 1.00
$\delta<1.25^{3}\uparrow$ Dyn 0.96 0.83 1.00 1.00 0.99
All 0.98 0.80 1.00 1.00 1.00
[…]
All 7.28E+05 5.69E+05 0.02 0.04 0.02
$\delta<1.25\uparrow$ Dyn 0.29 0.37 0.08 0.12 0.04
All 0.20 0.21 0.01 0.10 0.02
$\delta<1.25^{2}\uparrow$ Dyn 0.14 0.37 0.01 0.01 0.02
All 0.10 0.21 0.01 0.03 0.01
$\delta<1.25^{3}\uparrow$ Dyn 0.12 0.28 0.00 0.01 0.02
All 0.06 0.16 0.00 0.01 0.01
Table 8: Camera poses accuracy for out-of-domain data [58]. We show a comparison to previous methods on the predicted camera poses. We compare 3 ways of running our method. Ours (C): Our inference time outputs. Ours (C)+BA: Our inference time outputs, followed by a short Bundle Adjustment for a model that was trained only on cats. Ours (C) FT: The outputs of the model that was trained only on cats after fine-tuning our losses for each specific video
DROID-SLAM[44] ParticleSfM[61] RCVD[22] CasualSAM[59] Ours Ours Ours
(C) (C)+BA (C)+FT
Balloon1 ATE(mm) 2.87 5.81 1.7E+02 5.57 21.47 4.17 4.14
RPE T.(mm) 4.58 6.51 2.5E+02 4.96 27.73 6.76 6.76
RPE R.(deg.) 0.05 0.08 3.35 0.05 0.40 0.07 0.07
Balloon2 ATE(mm) 7.81 13.51 3.5E+02 7.74 41.25 10.22 9.92
RPE T.(mm) 12.82 14.16 3.9E+02 11.47 77.12 16.80 16.78
RPE R.(deg.) 0.13 0.13 3.21 0.11 0.84 0.17 0.17
DynamicFace ATE(mm) 2.80 9.71 1.1E+02 3.59 32.79 4.11 3.73
RPE T.(mm) 1.71 7.93 2.4E+02 2.05 48.39 3.20 3.16
RPE R.(deg.) 0.04 0.17 3.34 0.05 1.04 0.07 0.06
Jumping ATE(mm) 7.65 13.31 2.8E+02 7.74 24.24 8.38 8.61
RPE T.(mm) 10.25 11.27 2.8E+02 8.69 36.34 11.35 12.13
RPE R.(deg.) 0.05 0.06 3.09 0.05 0.21 0.07 0.08
Playground ATE(mm) 7.62 85.47 1.1E+02 5.45 27.44 6.47 5.06
RPE T.(mm) 9.51 90.10 3.3E+02 7.68 40.28 8.00 6.42
RPE R.(deg.) 0.10 0.75 4.69 0.11 0.40 0.10 0.10
Skating ATE(mm) 7.21 19.37 78.24 7.28 27.57 9.21 8.88
RPE T.(mm) 8.64 24.76 3.2E+02 8.65 45.02 11.19 11.44
RPE R.(deg.) 0.04 0.15 3.91 0.05 0.24 0.07 0.07
Truck ATE(mm) 22.55 - 1.1E+02 17.70 42.47 19.49 17.53
RPE T.(mm) 31.68 - 3.6E+02 30.24 69.22 34.61 28.37
RPE R.(deg.) 0.06 - 2.77 0.05 0.26 0.05 0.05
Umbrella ATE(mm) 5.20 39.45 66.01 7.38 39.27 7.33 5.99
RPE T.(mm) 8.11 12.08 3.7E+02 7.01 39.85 6.98 8.03
RPE R.(deg.) 0.04 0.05 3.05 0.03 0.20 0.03 0.03
Mean ATE(mm) 7.96 26.66 1.6E+02 7.81 32.06 8.67 7.98
RPE T.(mm) 10.91 23.83 3.2E+02 10.09 47.99 12.36 11.64
RPE R.(deg.) 0.07 0.20 3.43 0.06 0.45 0.08 0.08
STD ATE(mm) 6.25 28.13 1.0E+02 4.25 8.11 4.89 4.50
RPE T.(mm) 9.06 29.82 55.63 8.60 16.82 9.86 7.95
RPE R.(deg.) 0.03 0.25 0.61 0.03 0.32 0.04 0.04